OpenStudy (anonymous):

Simplify the following expression. x − 6y + 2z − 5x − 3y + 8z

OpenStudy (anonymous):

@jazy can you please help, you don't have too!

OpenStudy (anonymous):

Of course! Combine like terms. Remember to keep the sign in front of each term. x − 6y + 2z − 5x − 3y + 8z Let's do them seperately: x - 5x = -4x -6y - 3y = -9y 2z + 8z = 10z So, your simplified equation is: -4x - 9y + 10z

OpenStudy (anonymous):

thanks!!!! i need help with another one but i need to draw it! @jazy

OpenStudy (anonymous):

welcome.. ok!

OpenStudy (anonymous):